OpenStudy (anonymous):

Please help me with some questions! Fan and Medal!! . A restaurant investigated the correlation between prices and the number of customers. The linear model below represents this situation, where x is the average meal price [in dollars], and y is the number of daily customers. Interpret the slope. y = 437 - 18x If the average meal price increases by $1, the number of daily customers decreases by 18. B.The average number of daily customers is 18. C. If the average meal price decreases by $1, the number of daily customers decreases by 18. D. The average price per meal is $18.

OpenStudy (clongoria2):

The slope would be 18 because this is the rate of change. Which is what the slope represents. With that in mind. Can you give the answer?

OpenStudy (anonymous):

^ is correct.

OpenStudy (anonymous):

Thanks! Can someone help me with more?

OpenStudy (anonymous):

Gordon recently started jogging. Last week he jogged for 10 minutes. Starting this week, he decided to add 7 minutes to his jog each week. Which of the following equations represents how long his jog will be after x additional weeks? A. y = 10x + 17 B. y = 7x + 17 C. y = 7x + 10 D. y = 10x + 7